OpenStudy (anonymous):

MEDAL please help me. i dont understand. "Find the sum, if it converges. (1/3)-(1/9)+(1/27)-..."

OpenStudy (dumbcow):

its a geometric series r = -1/3

OpenStudy (dumbcow):

\[\text{\sum} = \frac{a_1}{1-r}\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }-\frac{ 1 }{ 9}+\frac{ 1 }{ 27}\] common denominator

OpenStudy (anonymous):

\[\frac{ 1*9 }{ 3*9 }-\frac{ 1*3 }{ ?9*3 }+\frac{ 1 }{ 27}=\frac{ 9-3+1 }{ 27 }=\]